# Wireless Datasets for Aerial Networks

### v1

@@ XI-B Dataset Format
Each location folder within testbed contains the following core files:
  • power_log.txt: Table II shows a snippet from the power_log.txt file, where each row represents a RSS measurement captured by the UAV receiver. The first column contains precise timestamps (in microseconds), and the third column provides the measured RSS in dB. Although the raw file also includes a middle column for sample indexing, it is of no use for signal analysis.
  • quality_log.txt: Has the same format as Table II, but contains RSQ values instead of RSS.
@@
  • log.csv: Contains time-synchronized UAV navigation data, including GPS coordinates (latitude, longitude, altitude), speed, heading, and satellite metadata. A sample snippet of this file is illustrated in Fig. 26.
  • angles.mat: Provides the UAV’s orientation in terms of roll, pitch, and yaw.
TABLE II: Sample Format of power_log.txt
@@
Figure 26: Sample entries from log.csv showing UAV navigation and GPS metadata.
XI-C Representative Results
((a))
((b))
((a))
((b))
((c))
Figure 28: Representative signal and mobility characteristics from the AFAR dataset across different teams and environments [49].
The measurements in the AFAR dataset exhibit considerable variability resulting from differences in UAV trajectories, UGV placements, and environment-specific propagation effects. KPIs such as RSS, RSQ, and UAV motion parameters (e.g., speed and altitude) vary significantly across locations. For instance, the UAV speed and altitude profile for Team-309, who employed a fixed waypoint trajectory, are shown in Fig. 28(a). As the UAV approaches each waypoint, its speed increases, then decreases upon arrival, before accelerating again toward the next target. This cyclical speed pattern is characteristic of waypoint-based navigation. Additionally, the UAV maintains a relatively constant altitude of approximately 30 meters throughout the mission.
Fig. 28(b) presents RSS heatmaps overlaid on the UAV flight paths. The diamond marker indicates the RF source location. As expected, RSS is strongest when the UAV is in close proximity to the source and weakens with increasing distance. However, the spatial distribution of received power is not uniform, reflecting the influence of multipath and shadowing effects.
To highlight the contrast between simulated and real-world signal behavior, Fig. 28(c) plots RSS and distance against time for both DT and real-world environments. In the DT environment, RSS remains relatively smooth and predictable. In contrast, the real-world data exhibits significant fluctuations, even at similar distances, due to dynamic factors such as fading, body blockage, and environmental clutter. This discrepancy underscores the importance of accounting for real-world propagation effects when developing and validating RF localization algorithms.
XI-D Possible Uses of Dataset
XI-D1 A2G Channel Propagation Modeling
The dataset enables realistic modeling of A2G wireless channels, accounting for UAV-specific factors such as altitude, elevation/azimuth angles, velocity, and orientation (roll, pitch, yaw). These measurements help characterize the propagation environment under mobility and elevation diversity [50].
XI-D2 Antenna Gain and Shadowing Analysis
The dataset facilitates the evaluation of directional antenna performance and gain variations due to UAV-body shadowing. This is particularly relevant for understanding signal attenuation in NLoS conditions and UAV maneuvers.
XI-D3 Performance Benchmarking
Researchers can perform comparative analysis of UAV-assisted RF localization algorithms. The dataset allows assessment of trajectory efficiency, signal quality, and overall localization accuracy under controlled and real-world constraints.
XI-D4 Data-Driven Localization Algorithms
The time-synchronized RF and positional data provide a robust basis for training and evaluating machine learning models for RF source localization, including regression, classification, or hybrid approaches.
XI-D5 Signal Strength and Fading Prediction
Deep learning models can be trained to forecast RSS/RSQ values under varying mobility conditions. This supports proactive planning in UAV-assisted sensing and communication tasks.
XI-D6 Flight Path Optimization
Using the dataset, UAV trajectories can be optimized for better link reliability, minimal energy use, or improved localization precision, either via reinforcement learning or optimization-based techniques.
XI-D7 Simulation-to-Reality Transfer Learning
The paired DT and real-world measurements enable transfer learning strategies that improve model robustness across synthetic and physical environments.
XII UAV Signal Classification Dataset
Figure 29: The experimental setup for detecting and capturing RF signals from UAV controllers. The input of the receiver has an LAN, a bandpass filter and a parabolic antenna receiver [55].
((a))
((b))
((c))
((d))
((e))
((f))
((g))
((h))
((i))
((j))
((k))
((l))
Figure 30: This shows a sample of RF signal captured from eight different UAV controllers and four different UAVs while on flight: (a) Graupner MC-32, (b) Spektrum DX6e, (c) Futaba T8FG, (d) DJI Phantom 4 Pro, (e) DJI Inspire 1 Pro, (f) JR X9303, (g) Jeti Duplex DC-16, (h) FlySky FS-T6, (i) DJI Matrice 600 UAV, (j) DJI Phantom 4 Pro UAV, (k) DJI Inspire 1 Pro UAV, (l) DJI Mavic Pro [55].
XII-A Description of Hardware and Software
The experimental setup is shown in Fig. 29. In the figure, a 24 dBi parabolic antenna operating in the 2.4 GHz ISM band listens to RF signals from UAV controllers, which are used to send control and navigation commands to an incoming UAV. The choice of the parabolic antenna is to further increase the range of the receiver because of the improved directivity (gain) of the antenna. The output signal from the antenna is fed to the receiver through an RF chain that combines a LNA and an RF bandpass filter. The LNA is used to amplify weak signals, while the bandpass filter ensures the input signal is band-limited so as not to saturate the receiver. Receiver saturation will cause non-linearity effects like intermodulation distortions (IMD), gain compression, and loss of signal integrity after amplification. For the experiment, the receiver is a high-resolution mixed signal Keysight oscilloscope (6 GHz Keysight MSOS604A) capable of sampling the captured signal at 20 GSa/s. This high sampling rate ensures the detection system captures all the transient features or fingerprints of the captured UAV RF controller signals. Also, within the receiver in Fig. 29, there is a custom MATLAB script for signal detection, data preprocessing and transformation of the raw signal.
Fig. 30 shows samples of UAV RF remote controller (RC) signals and UAV emitted signals captured using the detection systems in Fig. 29  [55]. From this figure, it is obvious that each of the UAV RF signal have some unique features/waveform that can be exploited by an intelligent signal recognition system to identify the UAV or its controller. Moreover, using advanced signal localization algorithms with multiple RF receivers, the UAV and its controller can be localized and monitored.
@@ XII-A Description of Hardware and Software
TABLE III: Drone catalogue for the classification signal dataset.
@@
TABLE IV: Metadata of UAS signal classification dataset.
@@
XII-B Dataset Format
The UAS signal classification dataset consists of RF signals from 17 popular commercial UAV controllers (15 of which are unique) from eight different manufacturers [54]. The UAV controller catalogue includes DJI, Futaba, Hobby King, and Turnigy as shown in Table III. For each UAV RF signal category in the dataset, Table IV provides basic information or metadata of each UAV RF samples in the UAS signal classification dataset. To extract and visualize a specific UAV RF signal and its metadata from the dataset, we need to create a MATLAB object instance, as in object-oriented programming. Also, to perform any experiment using the dataset, we can extract all or part of the dataset by creating a custom database.
@@ XII-B Dataset Format
The dataset is organized into processed CSV files for wireless network analysis and comprises three primary directories:
  1. 1.
    Creating a UAV Object to Visualize RF signals: To extract UAV RF signal metadata from a .mat file in the dataset, we use a MATLAB script (droneRC.m) that is provided with the UAS signal classification dataset. The droneRC.m script defines a MATLAB class, and several methods/functions are defined for creating a drone RF object (specific UAV RF signal) as shown in Fig. 31.
  2. 2.
    Creating a Database for Experiment: To create a database of UAV RF signal from the UAS signal classification dataset, we use the createDatabase function. This function can generate a database of drone controller RF signals in both matrix and table format. Using the function, you can specify the number of signals, features to include, and the format for the database as shown in Fig. 32.
((a))
((b))
Figure 33: Representation Result for UAV Detection and Classification Using the UAS signal classification dataset: (a) The performance of the proposed detection systems as a function of the SNR, the detection threshold ($\delta$), and the fixed FAR, (b) The performance of the ML-based model as a function of the SNR, the number of UAV RF controller signal category extracted from the UAS signal classification dataset [55].
XII-C Representative Results
From Fig. 33(a), we see that for a fixed false alarm rate (FAR), increasing the detection threshold ($\delta$) will reduce the detection accuracy. Also, from Fig. 33(b), we see that the accuracy of the classification system depends on the machine learning model deployed, the SNR, and the number of UAV classes in the database. Once again, the study shows that it is possible to achieve an accuracy of 98.13% in classifying 15 different UAV controllers using classical machine learning models like kNN and random forest. In [57], the authors investigated the impact of using the convolutional neural network (CNN) to classify/identify UAV RF controller signals. The CNN models are trained using spectrogram images representation of the raw UAV controller RF signals. The CNN model achieved an accuracy of about 92%.
XII-D Possible Uses of Dataset
XII-D1 Benchmarking New Detection and Classification Algorithms
As researchers continue to work in this domain, it is important to benchmark or compare the performance of their algorithms with alternative classical models in literature. In such a study, it is important to ensure the UAV RF data used were reliably captured in the same environment. Moreover, the performance of different algorithms as a function of thermal noise or antenna temperature can be studied comparatively.
XII-D2 Investigating the Impact of Intentional or Non-intentional Radiators
Researchers can use the UAS signal classification dataset to investigate the impact of intentional interferences like jamming signal on the performance of UAV RF-based detection and classification algorithms. This is important because of the array of new and existing UAV counter-measure EW systems. Also, as described in [55], researchers can use the UAS signal classification dataset to investigate the impact of in-band and out-of-band non-intentional interference signals like Wi-Fi, BLE, Zigbee, medical equipment, micro-wave oven, and others. This is important for electromagnetic compliance (EMC) certification especially in shared unlicensed spectrum.
XII-D3 Generating Synthetic UAV RF Dataset
Using the UAS signal classification dataset, researchers can generate synthetic UAV RF signals. For example, using generative adversarial networks (GANs) or variational autoencoders (VAEs), we can create large synthetic data using the UAS signal classification dataset as input. The synthetic data can be used to train more robust machine learning models, increasing the generalization and accuracy of such models. Other methods that can be used to create synthetic UAV controller RF data include data augmentation techniques (translation and rotation), kernel density estimation, parametric statistical models, random sampling and Monte Carlo simulation. As a first principle, all these methods will need an input dataset like the UAS signal classification dataset for training.
UAV flight path data, including throughput and RSRP readings, are crucial to developing, experimenting with, and evaluating next-generation wireless networks and DT solutions. Our datasets capture critical A2G propagation and link characteristics based on trajectory, as well as interactions with emulated and simulated environments. The advantage of the emulation environment is that it implements full-stack software as it runs in real-world SDRs (such as USRPs) and UAVs; on the other hand, it can be relatively slow and computationally intensive for development purposes. Simulation abstractions offer the advantage of accelerating the development of UAS and radio algorithms that can subsequently be evaluated in emulation and real-world testbeds, though this comes at the expense of reduced realism.
XIII-A Description of Hardware and Software
The measurements were obtained through a combination of MATLAB simulation and DT emulation [60], utilizing AERPAW’s cutting-edge experimental facilities at Lake Wheeler. The major components are as follows.
The UAV testbed includes an LTE SISO radio-equipped UAV, GPS receivers, and flight controllers that ensure accurate trajectory tracking. Field experiments were conducted in the Lake Wheeler area, where high-fidelity RSRP measurements were collected from four BSs (LW1–LW4). A MATLAB-based simulation environment was developed to model UAV flights in virtual settings, with field measurements used to validate the simulations under realistic conditions. The AERPAW DT further supports emulation of UAV-to-base-station communications, providing parameterized control over mobility patterns, radio environments, and network conditions.
There are custom MATLAB and Python scripts for data processing, analysis, and visualization. The full dataset, including all processing and simulation code, is made publicly available at [58, 59] to facilitate reuse as easily and transparently as possible.
XIII-B Dataset Format
The dataset is organized into processed CSV files for wireless network analysis and comprises three primary directories:
  1. 1.
@@ XIII-B Dataset Format
    Emulated RSRP Measurement Using Fixed Trajectory: Emulated data for the UAV flights on designed trajectories are found in the fixed_trajectory_rsrp_emulation folder. The dataset contains timestamps, altitude, latitude, longitude, and RSRP measurements, along with pitch, roll, and yaw.
  2. 2.
    Simulated RSRP Measurement Using Fixed Trajectory: Simulated data for the UAV flights on designed trajectories are found in the fixed_trajectory_rsrp_simulation folder. The dataset contains timestamps, altitude, latitude, longitude, and RSRP measurements.
  3. 3.
    Simulated RSRP Measurement Using Autonomous Trajectory: Simulated data for the UAV flights on autonomous trajectory are found in the autonomous_trajectory_rsrp_simulat ion folder. The dataset contains timestamps, altitude, latitude, longitude, and RSRP measurements.
Table V provides the structure of the dataset used in this section. Each row of the table corresponds to a specific timestamp in a UAV flight and records the UAV’s position (altitude, longitude, and latitude), RSRP, SNR, and data rate.
@@
TABLE V: UAV measurement dataset sample with RSRP, SNR, and data rate.
Time
Lon.
Lat.
Alt.
RSRP
SNR
Rate
@@
((a))
((b))
((c))
((d))
Figure 34: RSRP measurement with respect to (a) LW1, (b) LW2, (c) LW3, and (d) LW4 for fixed trajectory in emulation [60].
((a))
((b))
((c))
((d))
Figure 35: RSRP measurement with respect to (a) LW1, (b) LW2, (c) LW3, and (d) LW4 for autonomous trajectory in simulation [60].
To enable effective visualization and analysis, the dataset includes a collection of post-processing scripts. For scenarios with fixed UAV trajectories, scripts such as plot_rsrp_emulation.m and plot_rsrp_distance_emulation.m are used to graphically represent RSRP measurements as a function of UAV location and distance from a specified BS, respectively. The same set of plotting conventions is followed for the simulated dataset as well, with names or identifiers of scripts distinguishing simulation and emulation results. For the autonomous path scenario, plot_rsrp_simulation.m generates RSRP plots in dynamic UAV trajectories, and the Python scripts plot_throughput_distance_vs_time_lw34
.py and plot_throughput_distance_vs_time _lw12.py provide distance vs. time plots and throughput for each BS pair, respectively. These post-processing scripts streamline the workflow, enabling reproducible preparation of all key figures and statistical analysis presented in this paper.
XIII-C Representative Results
This part discusses analysis and results based on the dataset, particularly in terms of comparing simulation with emulation and the effect of trajectory design:
RSRP for Individual BSs: Here, we present the emulation RSRP measurements for the fixed trajectory in Fig. 34. When the UAV is closer to a BS, the signal strength tends to be stronger; e.g., we observe a strong signal strength in Fig. 34(a) because the UAV is close to the LW1 compared to other towers. Results in simulated environment (not provided here due to space reasons) show a high similarity in signal strength with the emulation results.
Fig. 35 shows the RSRP measurement by UAV to individual BSs for autonomous trajectory in simulation. We see a strong RSRP near all BSs except LW4 because UAV goes near LW1, LW2, and LW3, but LW4 is far away, and the UAV is restricted to the geofence. The trajectory details (fixed and autonomous) and other results, such as distance versus RSRP measurements and throughput measurements, are presented in  [60].
XIII-D Possible Uses of Dataset
The AERPAW Lake Wheeler UAV dataset is a useful tool for a range of wireless networking and DT research tasks:
XIII-D1 Simulation Environment Calibration
Use emulation data to calibrate, validate, and tune simulation environments, enabling them to be more effective predictive tools for real-world deployments.
XIII-D2 A2G Signal Propagation Analysis
Investigate altitude, trajectory, and environmental effects on RSRP and throughput for UAV-ground links.
XIII-D3 DT Performance Evaluation
Benchmark and validate the fidelity of network emulators of DTs by direct comparison with field data.
XIII-D4 Trajectory-Aware Algorithm Benchmarking
Develop and test path planning, handover, and resource allocation algorithms using realistic UAV trajectory and signal data.
XIII-D5 Empirical Propagation Model Fitting
Construct and calibrate A2G path loss models and spatial radio maps from measured RSRP and SNR data.
XIII-D6 Machine Learning for Predicting Link Quality
Train the classifiers or regressors to predict RSRP or throughput using spatial and mobility attributes.
XIII-D7 Coverage and Connectivity Mapping
Generate 2D/3D coverage maps for representing and assessing the radio network connectivity along complex UAV flight trajectories.
XIV Ray Tracing Simulation and Measurement Comparison Dataset
XIV-A Description of Hardware and Software
((a))
((b))
Figure 36: Trajectory and altitude of the signal coverage measurements and RT simulation (reused from [62]).
The predefined trajectory and altitude over time of the UAV are demonstrated in Fig. 36. The UAV takes off near the LW1 tower. After takeoff, the UAV sweeps LW2 and LW3 towers. The UAV returns to LW1 and lands on the ground. Each waypoint of the trajectory is highlighted with a letter, which corresponds to vertical lines in Fig. 36(b). The UAV and LW towers have SISO antenna setups with $3.3$ GHz carrier frequency. The RSS is recorded by a dual-channel USRP B210 and GNU Radio at each tower while the UAV transmits signals. The RSS is measured for $20$ ms for every $100$ ms intervals.
((a))
((b))
((c))
Figure 37: RSS measurement and RT simulation results with the predefined trajectory for the specific UAV locations A-F in Fig. 36 (reused from [62]).
XIV-B Dataset Format
The dataset and post-processing scripts are publicly available at [61]. The dataset consists of three MATLAB scripts and three data folders or .mat files, labeled with the prefixes “C” and “D”, respectively. The structure of the dataset and post-processing scripts can be summarized as follows.
  • D1_RSS_Measurements: This folder has $5$ subfolders RSS measurements in dB scale at $5$ different LW towers (LW1-LW5). The GPS information of the predefined trajectory of the UAV is also included in each subfolder.
@@ XIV-B Dataset Format
  • D2_RT_RSS_results_data.mat: This file includes the RT simulation results over the predefined trajectory of the UAV. The path coefficient and propagation delays are stored for each UAV position.
  • D3_RT_Measurement_RSS_postprocessed_da
@@
    ta.mat: This file contains the time duration of the measurement campaign, RSS results from RT simulation, and calibrated RSS results from the measurements.
  • C1_trajectory_altitude.m: This script plots the predefined trajectory and altitude over time as shown in Fig. 36.
@@
  • C2_RSS_postprocessing.m: This script is for post-processing and calibration of RT and measured data. This script is optional for reproducing the results because the output of this script, which is D3_RT_Measurement_RSS_postprocessed_da
    ta.mat, is already included in the dataset.
  • C3_RSS_comparison.m: The post-processed RSS results from RT and the measurement campaign can be plotted by this script, which is demonstrated in Fig. 37. The output of this script shows the $5$ figures of results for each tower.
XIV-C Representative Results
The RT simulation results of RSS and measurement from each tower are shown in Fig. 37, where $\mathrm{Z}$ indicates out-of-coverage area. Here, the measurements from different antennas of the dual-channel USRP at each tower are labeled as USRP1 and USRP2, respectively. The RT simulation results are consistent with the measurements at all tower cases. Fluctuations are observed in the measurement over the trajectory due to changes in the direction (roll/yaw/pitch) of the UAV at each waypoint, fading effects, and other factors. It is also worthwhile to note that a $10$ dB variation between USRP1 and USRP2 is observed in the LW1 and LW2 cases, which are highlighted in the figure with arrows. This variation can be attributed to LoS blockage and channel conditions due to the antenna orientations of USRP1 and USRP2, facing the UAV. Moreover, altitude-dependent out-of-coverage areas can be found in the RT simulation for LW3 at lower altitudes during takeoff and landing (time interval before $100$ s and after $850$ s).
XIV-D Possible Uses of Dataset
XIV-D1 Benchmarking RT Algorithms
The dataset provides RSS measurements from the towers and simulated results from the NVIDIA Sionna RT. Thus, the dataset can be used for validation and benchmarking of different RT approaches by comparing the simulation results.
XIV-D2 RF Coverage Analysis and Trajectory Planning
Based on a rural area with dense foliage, the dataset enables the prediction and design of reliable RF coverage by comparing with the desired performance requirements, such as required SNR levels, etc.
XIV-D3 UAV Communication Link Analysis
The dataset captures fluctuation in RSS over the predefined trajectory, allowing for detailed analysis of link performance, e.g., throughput, under realistic propagation conditions.
XIV-D4 Statistical Propagation Channel Modeling
Since the RT results include propagation delay and path coefficient information, the dataset can be used for propagation channel characterization by statistically analyzing this information.
XV Concluding Remarks
Aerial wireless connectivity is becoming an essential enabler of next-generation communication systems, including 5G-Advanced and 6G. This paper presented a diverse collection of open and well-documented datasets from the NSF AERPAW testbed, covering various radio technologies such as 5G, Wi-Fi, and LoRa, and captured using UAVs, helikites, programmable SDR nodes, and commercial UE. We discussed the technical and regulatory challenges associated with developing programmable aerial wireless platforms to collect such datasets, including the integration of SDRs, real-time localization with centimeter-level precision, testbed-wide time synchronization, and compliance with FAA and FCC requirements. The resulting datasets offer high spatial and temporal resolution, supporting a wide range of research activities in wireless communications, signal processing, and machine learning. These datasets are curated according to FAIR principles and are intended to support the academic and industrial research community. They enable rigorous evaluation of propagation models, data-driven algorithm design, and performance benchmarking in altitude-varying environments. Future work will involve expanding the dataset scope to include cooperative and mobile scenarios, incorporating advanced networking features, and aligning with emerging standards to inform data-driven regulatory policy.

### v2

@@ A-H AFAR Challenge Dataset File Organization
Core files in each testbed location folder: Each location-specific folder within the testbed directory contains the following files:
  • power_log.txt: Contains RSS measurements captured by the UAV receiver. Each row includes a timestamp (microsecond resolution), a sample index, and the measured RSS in dB. The sample index column is not used for signal analysis.
  • quality_log.txt: Contains RSQ measurements with the same format as power_log.txt.
@@
  • log.csv: Time-synchronized UAV navigation data, including GPS coordinates (latitude, longitude, altitude), speed, heading, and satellite metadata.
  • angles.mat: UAV orientation data, including roll, pitch, and yaw angles.
Table III provides an example snippet of the power_log.txt file format, and Table IV illustrates representative entries from the log.csv file.
TABLE III: Sample Format of power_log.txt
@@
TABLE IV: Sample entries from log.csv showing UAV navigation and GPS metadata.
TimeUS Status GMS GWk NSats HDop Lat Lng Alt Spd GCRs VZ Yaw U
2166080919 5 3.27E+08 2292 14 0.77 35.727371 -78.6962127 112.65 0.03985 72.47443 0.007 0 1
2166280903 5 3.27E+08 2292 14 0.77 35.727371 -78.6962128 112.66 0.028284 261.8699 -0.059 0 1
2166460908 5 3.27E+08 2292 14 0.77 35.7273709 -78.6962131 112.66 0.04639 172.5686 0.007 0 1
2166660988 5 3.27E+08 2292 14 0.77 35.727371 -78.6962132 112.65 0.061294 5.61758 0.21 0 1
2166860968 5 3.27E+08 2292 14 0.77 35.727371 -78.6962131 112.66 0.023087 252.3499 -0.189 0 1
2167060890 5 3.27E+08 2292 14 0.77 35.7273709 -78.6962132 112.68 0.011664 329.0363 -0.148 0 1
2167280929 5 3.27E+08 2292 14 0.77 35.7273708 -78.6962132 112.69 0.008485 315 -0.13 0 1
A-I UAV Signal Classification Dataset Organization and Utilities
Dataset contents: The UAV signal classification dataset consists of processed RF signal recordings and associated metadata stored in MATLAB .mat files. The dataset includes RF signals from 17 commercial UAV controllers representing eight manufacturers, as summarized in Table V. Metadata for each RF sample, including controller identity and acquisition parameters, are summarized in Table VI.
@@ A-I UAV Signal Classification Dataset Organization and Utilities
TABLE V: UAV platforms and remote controllers included in the drone signal classification dataset.
@@
TABLE VI: Metadata of UAS signal classification dataset.
@@
Signal extraction and visualization: The dataset provides a MATLAB class definition, droneRC.m, which enables object-oriented access to individual UAV controller RF signals and their associated metadata. This class supports extraction, visualization, and preprocessing of RF samples stored in the dataset. Fig. 36 illustrates the use of the droneRC class for accessing RF signals and metadata.
Dataset construction for experiments: To support flexible experimental design, the dataset includes a MATLAB utility function, createDatabase.m, which enables construction of custom datasets from selected UAV controller signals. The function can generate databases in matrix or table format and allows users to specify the number of samples, selected controllers, and feature representations. Fig. 37 illustrates the database construction process.
These utilities enable reproducible benchmarking of UAV RF signal detection and classification algorithms under controlled experimental conditions, including evaluation across SNR levels, controller diversity, and interference scenarios.
@@
A-J Trajectory-Aware RSRP and Throughput Dataset File Organization
Directory organization: The dataset is organized into three primary directories corresponding to the measurement methodology and trajectory type:
@@ A-J Trajectory-Aware RSRP and Throughput Dataset File Organization
  • fixed_trajectory_rsrp_emulation: Emulated RSRP measurements collected along predefined UAV trajectories. Each record includes timestamps, UAV position (latitude, longitude, altitude), orientation (pitch, roll, yaw), and RSRP values.
  • fixed_trajectory_rsrp_simulation: Simulated RSRP measurements generated for the same fixed UAV trajectories. Each record includes timestamps, UAV position, and RSRP values.
  • autonomous_trajectory_rsrp_simulation: Simulated RSRP measurements collected along autonomous UAV trajectories subject to geofencing constraints. Each record includes timestamps, UAV position, and RSRP values.
CSV file structure: Each CSV file corresponds to a UAV flight and contains time-indexed measurements including UAV position (longitude, latitude, altitude), radio metrics such as RSRP and SNR, and application-layer throughput where applicable. Table VII illustrates a representative sample of the CSV structure.
Post-processing utilities: The dataset repository includes MATLAB and Python scripts used to generate the figures and analyses presented in the paper. These scripts support visualization of RSRP as a function of UAV location, distance to base stations, and time, as well as throughput analysis for fixed and autonomous trajectories.
These utilities enable reproducible preparation of all representative figures and statistical results reported in [37].
@@
TABLE VII: UAV measurement dataset sample with RSRP, SNR, and data rate.
Time (s) Lon. (deg) Lat. (deg) Alt. (m) RSRP (dBm) SNR (dB) Rate (Mbps)
@@
A-K Ray-Tracing Simulation and Measurement Comparison Dataset File Organization
The dataset is distributed as a collection of data folders and MATLAB scripts labeled using the prefixes “D” (data) and “C” (code). The organization supports reproducible comparison between measured RSS and ray-tracing simulation results.
Data components:
  • D1_RSS_Measurements: Contains subfolders corresponding to RSS measurements (in dB) collected at five base station locations (LW1–LW5). Each subfolder also includes the GPS information associated with the predefined UAV trajectory.
@@ A-K Ray-Tracing Simulation and Measurement Comparison Dataset File Organization
  • D2_RT_RSS_results_data.mat: Stores RT simulation outputs for the predefined UAV trajectory, including path coefficients and propagation delays at each UAV position.
@@
  • D3_RT_Measurement_RSS_postprocessed_data.mat: Contains the time duration of the measurement campaign, RSS values derived from RT simulation, and calibrated RSS values obtained from measurements.
Post-processing scripts:
  • C1_trajectory_altitude.m: Plots the predefined UAV trajectory and altitude as a function of time.
@@
  • C2_RSS_postprocessing.m: Performs post-processing and calibration of RT and measured RSS data. This step is optional for reproducing the published results, as the calibrated outputs are included in D3_RT_Measurement_RSS_postprocessed_data.mat.
  • C3_RSS_comparison.m: Generates comparative plots of measured and RT-simulated RSS values for each base station, corresponding to the results presented in the main text.
These components enable full reproduction of the reported ray-tracing comparison results.